8 years ago the father's age was 8 times the age of his son.after 10 years,the father's age will be twice the age of the son.what are their present age
Answer by lwsshak3(11628)

8 years ago the father's age was 8 times the age of his son.after 10 years,the father's age will be twice the age of the son.what are their present age
***
8 yrs ago:
let x= son's age
8x=father's age
At present:
x+8=son's age
8x+8=father's age
After 10 years:
x+18=son's age
8x+18=father's age
father's age twice son's age
8x+18=2(x+18)
8x+18=2x+36
6x=18
x=3
Present ages:
x+8=11=son's age
8x+8=32=father's age